An ideal Stirling engine using helium as the working fluid operates between temperature limits of 300 and 2000 K and pressure limits of 150 kPa and 3 MPa. Assuming the mass of the helium used in the cycle is 0.12 kg, determine (a) the thermal efficiency of the cycle, (b) the amount of heat transfer in the regenerator, and (c) the work output per cycle.


1
Expert's answer
2021-11-29T15:13:59-0500

Draw the T-s diagram of the Stirling cycle as shown below.



(a)

Obtain the characteristic gas constant and specific heat at constant volume for helium gas at 300 K from the table.

Characteristic gas constant,

R = 2.0769 kJ/kgK

Specific heat at constant volume,

Cv = 3.1156 kJ/kgK

Calculate the thermal efficiency of ideal Stirling cycle.

"η_{th} = 1 -\frac{T_L}{T_H}"

Here, TL is the minimum temperature of the cycle and TH is the maximum temperature of the cycle.

Substitute 300 K for TL and 2000 K for TH.

"η_{th} = 1 -\frac{300}{2000} \\

= 0.85 \\

= 85 \; \%"

Therefore, the thermal efficiency of the ideal Stirling cycle is 85 %.

(b)

Calculate the amount of heat transferred in the regenerator.

"Q_{regen} = mC_v(T_1-T_4)"

Here, m is the mass of helium used in the cycle, Cv is the specific heat at constant volume of helium, T1 is the temperature of the helium gas at the exit of the regenerator (maximum cycle temperature) and T4 is the temperature of the helium gas at the inlet of the regenerator (minimum cycle temperature).

Substitute 0.12 kg for m, 3.1156 kJ/kgK for Cv, 300 K for T4 and 2000 K for T1.

"Q_{regen} = 0.12 \times 3.1156(2000-300) \\ 

= 635.58 \; kJ"

Therefore, the amount of heat transferred in the regenerator is 635.58 kJ.

(c)

Obtain the relation for the ratio of specific volumes at the end and beginning of the isothermal heat addition using the combined gas law as follows

"\frac{P_1v_1}{T_1} = \frac{P_3v_3}{T_3} \\

 \frac{v_3}{v_1} = \frac{T_3P_1}{T_1P_3}"

Here, v1 is the specific volume of helium at the beginning of isothermal heat addition, v3 is the specific volume of helium at the end of isochoric heat rejection process, P1 is the pressure of helium at the beginning of isothermal heat addition, P3 is the pressure of helium at the end of isochoric heat rejection process, T1 is the temperature of helium at the beginning of isothermal heat addition and T3 is the temperature of helium at the end of isochoric heat rejection process,

The process 2 to 3 is at constant volume. Therefore, the above relation becomes.

"\frac{v_2}{v_1} = \frac{T_3P_1}{T_1P_3}"

Here, v1 is the specific volume of helium at the beginning of isothermal heat addition, v2 is the specific volume of helium at the end of isothermal heat addition, P1 is the pressure of helium at the beginning of isothermal heat addition, P2 is the pressure of helium at the end of isochoric heat rejection process, T1 is the temperature of helium at the beginning of isothermal heat addition and T2 is the temperature of helium at the end of isochoric heat rejection process,

Substitute 2000 K for T1, 300 K for T3, 3000 kPa for P3 and 150 kPa for P1.

"\frac{v_2}{v_1} = \frac{300 \times 3000}{2000 \times 150} =3"

Calculate the change in entropy during the isothermal heat addition.

"s_2-s_1 = R ln(\frac{v_2}{v_1})"

Here, v1 is the specific volume of helium at the beginning of isothermal heat addition, v2 is the specific volume of helium at the end of isothermal heat addition and R is the characteristic gas constant of helium.

Substitute 2.0769 for R, 3 for "\frac{v_2}{v_1}"

"s_2-s_1 = 2.0769 ln(3) = 2.2817 \;kJ/kg \cdot K"

Calculate the heat added during the isothermal heat addition process.

"Q_{in} = mT_1(s_2 -s_1)"

Here, m is the mass of helium used in the cycle, T1 is the temperature of helium at the beginning of isothermal heat addition and (s2 -s1) is the change in entropy during the isothermal heat addition.

Substitute 0.12 kg for m, 2000 K for T1 and 2.2817 kJ/kgK for (s2-s1).

"Q_{in} = 0.12 \times 2000 \times 2.2817 \\

= 547.608 \; kJ"

Calculate the net-work output per cycle.

"W_{net} = \u03b7_{th}Q_{in}"

Here, Qin is the heat added during the isothermal heat addition process and ηth is the thermal efficiency of the ideal Stirling cycle.

Substitute 0.85 for ηth and 547.608 kJ for Qin.

"W_{net} = 0.85 \times 547.608 \\

= 465.467 \; kJ"

Therefore, the net-work output per cycle is 465.467 kJ.